# Clue-Guided Path Exploration: An Efficient Knowledge Base Question-Answering Framework with Low Computational Resource Consumption

### v1

@@ 4.2 Question-Answering
We conducted experiments on three question-answering sets: 1-hop Judge set, multi-hop Judge set, and 1-hop Query. In the Judge set, the knowledge chains corresponding to questions with negative answers are incomplete in the knowledge base. We believe that accurate path exploration should be able to discover incomplete knowledge chains. However, for incomplete knowledge chains, TOG tends to discard them and requires LLMs to directly answer the questions. To facilitate a fair comparison, we modified this logic and designated it as TOG*.
@@
The results are shown in Table 1. It is evident that the baseline methods, IO and CoT, which do not utilize external knowledge, exhibit very low accuracy rates in answering questions. This indicates that MOOC Q&A demands a high reliance on the knowledge base, making it well-suited for evaluating the effectiveness of knowledge base question-answering methods.
@@
On the Judge sets, CGPE demonstrates outstanding performance, surpassing the previous state-of-the-art method TOG. Even the performance of CGPE based on ChatGLM3-6B surpasses that of TOG based on GPT-4. If the answering stage model is switched to ChatGPT, while the other stages use ChatGLM3-6B, CGPE shows a substantial improvement, approaching the effectiveness of methods based on GPT-4. This suggests that, for the Judge set, our proposed CGPE has achieved impressive results in path exploration on ChatGLM3-6B, albeit constrained by the answering capability of ChatGLM3-6B.
@@
On the Query set, considering the partial match metric, TOG based on ChatGPT and GPT-4 performs reasonably well, while its effectiveness diminishes on ChatGLM3-6B. In contrast, our proposed CGPE framework outperforms TOG on ChatGLM3-6B, delivering comparable results to TOG when based on ChatGPT and GPT-4. Examining the exact match metric, CGPE consistently outperforms TOG across all scenarios. This indicates that while TOG can identify partial answers to questions, it struggles to comprehensively retrieve all possible answers to a given question. In contrast, CGPE excels in providing a comprehensive set of answers to the posed questions. Even when substituting the answering-stage model with ChatGPT and utilizing ChatGLM3-6B for other stages, CGPE demonstrates substantial improvement. This underscores the applicability of our proposed method to low-parameter models and emphasizes its practical value.
4.3 Computational Cost
@@
In the knowledge base question-answering process, the majority of computational resources are dedicated to utilizing the Large Language Model, while the remaining computational consumption is relatively negligible. We calculate the average number of LLM calls during the question-answering process to illustrate the computational resource consumption of TOG and CGPE. The results are shown in Table 2. In most cases, TOG has a significantly higher frequency of calls to LLMs, several times more than the calls made by CGPE to LLMs. Additionally, regardless of whether it is dealing with single-hop or multi-hop problems, the frequency of TOG calls remains relatively stable. This is primarily because TOG does not have a clear termination condition; instead, it requires LLMs to determine whether termination is appropriate.
@@
Table 3: Ablation Studies of Matching Step on 1-Hop Judge set. The steps for answering questions are based on ChatGPT, while all other steps are performed using ChatGLM3-6B.
4.4 Ablation Studies
4.4.1 Ablation Study on Path Exploration
CGPE comprises three key stages: clue extraction, path exploration, and answering. Clue extraction serves the path exploration stage, and it cannot be experimented as an independent stage. For answering stage, we adopted a straightforward approach. Thus, we only conducted ablation studies focusing on the path exploration stage. The path exploration stage can be divided into three parts: character matching, large language model pruning, and termination of exploration. The large language model pruning includes fuzzy matching, rough matching and exact matching. The results of ablation studies are shown in Table 3.
We treat direct questioning of ChatGPT as the blank control group, progressively incorporating LLM Pruning, character matching, and termination of exploration to validate the effects of each step. When not using the termination of exploration, we explored each path thrice, obtaining knowledge chains with three nodes. In these experiments, we utilized ChatGPT for answering questions, while all other steps employed ChatGLM. From the table 3, it is evident that solely employing LLM pruning step yields most of the benefits of the complete process, but the invocation of the large language model remains high at 8.98 times.
@@ 4.4.1 Ablation Study on Path Exploration
As a summary, the clue-guided LLM pruning step constitutes the majority of the overall stage effectiveness, while the roles of the other two steps primarily involve reducing computational costs.
@@
G4 G4 G4 80.5 71.4
Table 4: Performances with different backbone models for each stage. PE represents the path exploration stage, CE represents the clue extraction stage, and A represents the answering stage. GLM represents ChatGLM3-6B, CG represents ChatGPT, and G4 represents GPT-4.
4.4.2 Performances With Different Backbone Models
@@
Thanks to the plug-and-play nature of CGPE, we conducted experiments on the three stages of the framework—Clue Extraction, Path Exploration, and Answering—using different large language models with varying capabilities: ChatGLM3-6B, ChatGPT, and GPT-4. The purpose of these experiments is to validate the impact of models with varying capabilities on the effectiveness at each stage. These experiments were conducted on the Query set because, as demonstrated in Table 1, different models have a significant impact on the experimental results. It is conducive to showcasing the diverse influences of various models on each stage.
@@
The results are shown in Table 4. It is evident that substituting a model with higher capabilities at any stage enhances the final result, albeit with varying degrees of improvement across stages. During the answering stage, transitioning from ChatGLM to ChatGPT leads to noticeable improvements. However, the incremental gain is marginal when upgrading from ChatGPT to the more advanced GPT-4, as evidenced by the results on Judge sets in Table 1. This observation suggests that our framework may have a threshold in terms of the model’s capabilities required for the answering stage, where ChatGPT already demonstrates substantial adequacy. In the stages of clue extraction and path exploration, substituting a model with higher capabilities corresponds to superior results, with the most substantial enhancement observed in the path exploration stage. The path exploration stage primarily demands capabilities in information matching. If we fine-tune a less parameterized large language model specifically for “information matching” task Hu et al. (2021), it might achieve satisfactory results. Notably, clue extraction and answering require only one invocation of a large language model each, with the majority of computational costs concentrated in the path exploration stage. A less parameterized large language model for path exploration has the potential to significantly reduce the computational costs of CGPE during actual deployment.
5 Conclusion
@@
In this paper, we proposed a clue-guided path exploration framework named CGPE, designed to assist large language models in answering questions by exploring knowledge paths from knowledge bases. For a given question, CGPE initially extracts clues from the question and then explores knowledge paths guided by these clues. At each node along the knowledge path, we require a match with the extracted clues. Upon matching all clues within a knowledge path, we terminate the exploration of that path and present it to LLMs for answering the question. Experiments conducted on various question-answering sets indicate that CGPE outperforms previous state-of-the-art methods. Moreover, CGPE demonstrates high applicability to low-parameter LLMs compared to existing approaches, making it particularly valuable in scenarios with limited computational resources. Ablation studies validate the effectiveness of each step in the CGPE framework, providing insights for subsequent optimizations. In our future work, we aim to enhance the information matching capability of low-parameter LLMs to achieve better performance with lower computational resource consumption in CGPE.

### v2

@@ Introduction
We conducted experiments on three question-answering sets of MOOC Q&A: 1-hop Judge set, multi-hop Judge set, and 1-hop Query. In the Judge set, the knowledge paths corresponding to questions with negative answers are incomplete in the knowledge base. However, for incomplete knowledge paths, ToG tends to discard them and requires LLMs to directly answer the questions. To facilitate a fair comparison, we modified this logic and designated it as ToG*.
@@
The results are shown in 1. It is evident that the baseline methods, IO and CoT, which do not utilize external knowledge, exhibit very low accuracy rates in answering questions. This implies that these questions are not familiar to LLMs, and answering them requires a high degree of reliance on knowledge within the knowledge graph.
@@
On the Judge sets, CGPE demonstrates outstanding performance, surpassing the previous state-of-the-art method ToG. Even the performance of CGPE based on ChatGLM3-6B surpasses that of ToG based on GPT-4. If the answering stage model is switched to ChatGPT, while the path exploration stage uses ChatGLM3-6B, CGPE shows a substantial improvement, approaching the effectiveness of methods based on GPT-4. This suggests that, for the Judge set, our proposed CGPE has achieved impressive results in path exploration on ChatGLM3-6B, albeit constrained by the answering capability of ChatGLM3-6B. This fully aligns with our expectations: CGPE is specifically designed for path exploration in low-computing resource scenarios to enhance performance.
@@
On the Query set, considering the partial match metric, ToG based on ChatGPT and GPT-4 performs reasonably well, while its effectiveness diminishes on ChatGLM3-6B. In contrast, our proposed CGPE framework outperforms ToG on ChatGLM3-6B, delivering comparable results to ToG when based on ChatGPT and GPT-4. Examining the exact match metric, CGPE consistently outperforms ToG across all scenarios. This indicates that while ToG can identify partial answers to questions, it struggles to comprehensively retrieve all possible answers to a given question. In contrast, CGPE excels in providing a comprehensive set of answers to the posed questions.
The results of WebQuestions and WebQSP are shown in 2. Due to economic considerations and OpenAI’s policies, we did not test ChatGPT and GPT-4 for WebQSP and WebQuestions. From the baseline, we can observe that LLMs can answer a considerable portion of the questions solely based on their own knowledge, indicating that these questions are relatively familiar to LLMs. To more clearly demonstrate the ability of ToG and CGPE to retrieve knowledge from the knowledge graph, we require LLMs to answer questions based only on the retrieved knowledge, resulting in the outcomes of ToG - IO and CGPE - IO respectively. However, even with this requirement, the final results may still be relatively inflated due to the influence of the LLMs’ inherent knowledge. It can be seen that for relatively familiar questions, the results of CGPE are still significantly higher than those of ToG. This again demonstrates the effectiveness of our proposed method. The CGPE - IO results are better than those of ToG - IO, but not as much as the advantage of CGPE over ToG. This suggests that the knowledge retrieved by CGPE - IO is more unfamiliar to LLMs, while the knowledge retrieved by ToG - IO is more familiar to LLMs, which also reflects the effectiveness of our method for unfamiliar questions.
Computational Cost
@@
In the knowledge base question-answering process, the majority of computational resources are dedicated to utilizing the Large Language Model, while the remaining computational consumption is relatively negligible. We calculate the average number of LLM calls during the question-answering process to illustrate the computational resource consumption of ToG and CGPE. The results are shown in Table 3.
In most cases, ToG has a significantly higher frequency of calls to LLMs, several times more than the calls made by CGPE to LLMs. Additionally, regardless of whether it is dealing with single-hop or multi-hop problems, the frequency of ToG calls remains relatively stable. This is primarily because ToG does not have a clear termination condition; instead, it requires LLMs to determine whether termination is appropriate.
Table 3: Average Number of Calls to LLMs
@@
Table 4: Ablation Studies of Matching Step on 1-Hop Judge set. The answering step is based on ChatGPT, while all other steps are performed using ChatGLM3-6B.
@@
Ablation Studies
Ablation Study on Path Exploration
We conducted ablation studies on the path exploration stage, and the results of ablation studies are shown in Table 4.
We treat direct answering from ChatGPT as the baseline, progressively incorporating three step matching, character matching, and termination of exploration to validate the effects of each step. When not using the termination of exploration, we explored each path thrice, obtaining knowledge chains with three nodes. In these experiments, we utilized ChatGPT for answering questions, while all other steps employed ChatGLM3-6B. From the table 4, it is evident that solely employing three step matching yields most of the benefits of the complete process, but the invocation of the large language model remains high at 8.98 times.
@@
As a summary, the three step matching constitutes the majority of the overall stage effectiveness, while the roles of the other two steps primarily involve reducing computational costs.
Table 5: Performances with different backbone models for each stage. PE represents the path exploration stage, CE represents the clue extraction stage, and A represents the answering stage. GLM3 represents ChatGLM3-6B, CG represents ChatGPT.
@@
GPT-4 GPT-4 GPT-4 80.5 71.4
Performances With Different Backbone Models
@@
Thanks to the plug-and-play nature of CGPE, we conducted experiments on the three stages of the framework—Clue Extraction, Path Exploration, and Answering—using different large language models with varying capabilities: ChatGLM3-6B, ChatGPT, and GPT-4. The purpose of these experiments is to validate the impact of models with varying capabilities on the effectiveness at each stage.
@@
The results are shown in Table 5. It is evident that substituting a model with higher capabilities at any stage enhances the final result, albeit with varying degrees of improvement across stages. During the answering stage, transitioning from ChatGLM to ChatGPT leads to noticeable improvements. However, the incremental gain is marginal when upgrading from ChatGPT to the more advanced GPT-4, as evidenced by the results on Judge sets in Table 1. This observation suggests that ChatGPT is already capable of performing the task of “answering questions based on knowledge paths” very well. In the stages of clue extraction and path exploration, substituting a model with higher capabilities corresponds to superior results, with the most substantial enhancement observed in the path exploration stage. Notably, clue extraction and answering require only one invocation of a large language model each, with the majority of computational costs concentrated in the path exploration stage. Fine-tuning a less parameterized large language model specifically for “information matching” task (Hu et al. 2021) might achieve satisfactory results while reducing the computational costs.
Conclusion
@@
In this paper, we proposed a clue-guided path exploration framework named CGPE, designed to assist large language models in answering questions by exploring knowledge paths from knowledge bases. To address the information black box issue of existing methods, we discussed the composition of necessary information and designed a path exploration mechanism for CGPE centered around information. All steps of the mechanism are realized through a single information matching task, reducing the requirements for the comprehensive abilities of LLMs and enhancing the ability to solve unfamiliar problems. Experiments on multiple datasets have demonstrated the effectiveness of our design. CGPE not only outperforms existing methods in terms of results, but also requires far fewer calls to LLMs than existing methods. Moreover, CGPE demonstrates high applicability to low-parameter LLMs compared to existing approaches, making it particularly valuable in scenarios with limited computational resources. Ablation studies validate the effectiveness of each step in the CGPE framework, providing insights for subsequent optimizations. In our future work, we aim to enhance the information matching capability of low-parameter LLMs to achieve better performance with lower computational resource consumption in CGPE.